# Attacking LLM Watermarks by Exploiting Their Strengths

### v1

@@ Abstract
Advances in generative models have made it possible for AI-generated text, code, and images to mirror human-generated content in many applications. Watermarking, a technique that aims to embed information in the output of a model to verify its source, is useful for mitigating misuse of such AI-generated content. However, existing watermarking schemes remain surprisingly susceptible to attack. In particular, we show that desirable properties shared by existing LLM watermarking systems such as quality preservation, robustness, and public detection APIs can in turn make these systems vulnerable to various attacks. We rigorously study potential attacks in terms of common watermark design choices, and propose best practices and defenses for mitigation—establishing a set of practical guidelines for embedding and detection of LLM watermarks.
@@ 1 Introduction
Modern generative modeling systems have notably enhanced the quality of AI-produced content [BMR+20, SCS+22, Ope23a, Ope22]. For example, large language models (LLMs) like those powering ChatGPT [Ope22] can generate text that closely resembles human-crafted sentences. While this has led to exciting new applications of machine learning, there is also growing concern around the potential for misuse of these models, leading to a flurry of recent efforts on developing techniques to detect AI-generated content. A promising approach in this direction is to embed invisible watermarks into model-derived content, which can then be extracted and verified using a secret watermark key [KGW+23a, FGJ+23, CGZ23, KTHL23, ZALW24, KGW+23b, HCW+23, WHZH23, WYC+23].
Unfortunately, in this work, we identify that many of the key properties that make existing LLM watermarks successful also render them susceptible to attack. We study a number of simple, scalable attacks that can take advantage of common properties of existing watermarking schemes, including:
@@
    Robustness of the watermarks to potential modifications in the output text, so that the watermarks cannot be easily removed [KTHL23, KGW+23a, ZALW24, KTHL23, CGZ23, HCW+23];
@@
    Quality preservation of the output watermarked tokens, whose distribution is often desired to be close to the unwatermarked tokens to ensure the watermark does not degrade model utility [KGW+23a, FGJ+23, KTHL23, CGZ23, HCW+23, WHZH23]; and
@@
    Public detection APIs, which allow the general public to easily verify whether or not candidate text is AI-generated [FGJ+23, KGW+23a, SBC+19, MLK+23].
While these common properties of existing watermarking systems have clear benefits, we demonstrate that they can also enable malicious actors to easily launch two types of attacks: 1) watermark-removal attacks, which remove the watermark from the watermarked content, and 2) spoofing attacks, which create (potentially toxic) content with a target watermark embedded, thus making the content appear to be generated by a specific LLM when it is in fact not (see motivating examples in Table 1).
@@
Z-Score: 5.985.985.98 $\uparrow$
PPL: 4.464.464.46
(a) Spoofing attack Alan Turing was born in 1950 and died in 1994, at the age of 43. He was the brilliant American scientist and mathematician who is largely credited with being the father of modern computer science. He is known for his contributions to mathematical biology and musicology. He was also one of the pioneers of computer science...
@@
Exploiting quality-preserving
Z-Score: 2.402.402.40 $\downarrow$
PPL: 4.054.054.05
@@
Z-Score: 1.471.471.47 $\downarrow$
PPL: 4.574.574.57
Table 1: Examples generated using Llama-2-7B with/without the KGW watermark [KGW+23a] under various attacks. We mark the tokens in the green and red lists (see Appendix A). Z-score reflects the confidence of the watermark and perplexity (PPL) indicates the quality of the text. (a) In the spoofing attack, we exploit the robustness property of LLMs by generating incorrect content that appears as watermarked (matching the Z-score of the watermarked baseline), potentially damaging the reputation of the LLM. Incorrect tokens modified by the attacker are marked in orange and watermarked tokens in blue. (b-c) In watermark-removal attacks, attackers can effectively lower the Z-score below the detection threshold while preserving a high sentence quality (low PPL) by exploiting either the (b) quality-preserving property or (c) publicly available watermark detection API.
Our work rigorously explores removal and spoofing attacks for LLM watermarks. In doing so, we identify a number of critical trade-offs between watermark security and utility that emerge as a result of watermarking design choices. With these trade-offs in mind, we then propose potential defenses as well as a set of general guidelines to better enhance the security of next-generation LLM watermarking systems. Overall we make the following contributions:
  • We study how watermark robustness, despite being a desirable property to mitigate watermark removal attacks, can make the resulting systems highly susceptible to spoofing attacks, and show that challenges exist in detecting these attacks given that a single token can render an entire sentence inaccurate (Sec. 4).
  • We demonstrate that preserving quality of the output by using multiple watermarking keys can make the system susceptible to watermark-removal attacks (Sec. 5). We show both theoretically and empirically that a smaller number of keys may be necessary to defend against removal attacks, potentially at the cost of sacrificing output quality.
  • Finally, we identify that public watermark detection APIs can be exploited by attackers to launch both watermark-removal and spoofing attacks (Sec. 6). We propose a defense using techniques from differential privacy to effectively counteract spoofing attacks, and recommend setting query rate limits on the detection API and verifying the identity of users as simple mitigation strategies.
Throughout, we explore our attacks on three state-of-the-art watermarking mechanisms (KGW [KGW+23a], Unigram [ZALW24] and Exp [KTHL23]) and two LLMs (Llama-2-7B [TMS+23] and OPT-1.3B [ZRG+22])—demonstrating that these vulnerabilities are common to existing LLM watermarks, and providing caution for the field in deploying current solutions. Code is available at https://github.com/Qi-Pang/LLM-Watermark-Attacks.
@@ 2 Related Work
Advances in large language models (LLMs) have given rise to increasing concerns that such models may be misused for applications such as spreading misinformation, phishing, and academic cheating. In response, numerous recent works have proposed watermarking schemes as a tool for detecting LLM-generated text to mitigate potential misuse [KGW+23a, FGJ+23, CGZ23, KTHL23, ZALW24, KGW+23b, HCW+23, WHZH23, WYC+23]. These approaches involve embedding invisible watermarks into the model-generated content, which can then be extracted and verified using a secret watermark key.
Existing watermarking schemes share a few natural goals: (1) the watermark should be robust in that it cannot be easily removed; (2) the watermark should preserve the quality of the output text (usually quantified via unbiasedness in terms of the token distribution); and (3) the watermark should be easy to detect when given new candidate text. Unfortunately, we show that existing methods that aim to achieve these goals also render the resulting systems susceptible to watermark-removal or spoofing attacks.
Removal attacks. Several recent works have highlighted that paraphrasing methods may be used to evade the detection of AI-generated text [KSK+23, IWGZ18, LJSL18, LCW21, ZEF+23], with [KSK+23, ZEF+23] demonstrating effective watermark removal using a local LLM. However, these methods require expensive additional training for sentence paraphrasing which significantly impacts sentence quality, or assume a high-quality oracle model to guarantee the output quality is preserved. More fundamentally, our work differs in that we aim to directly connect and study how the inherent properties of watermarking schemes (such as quality preservation and detection APIs) can inform such removal attacks.
Spoofing attacks. [SKB+23, GLLH23] are the only works we are aware of that explore spoofing on watermarked LLMs. However, the work of [SKB+23] requires an unrealistic number of queries from attackers (1 million), limiting their method to one watermarking scheme [KGW+23a] and making it difficult to generalize to other watermarks. [GLLH23] train a new model to learn the watermarked token distribution, which can be infeasible for computation-bounded attackers, especially given the large number of queries to construct the training data and the cost of training a new LLM. The spoofing attacks we propose in the public detection setting can be generalized across all types of watermarks while requiring only a handful of queries to identify each token. Additionally, we are the first to explore the natural tension that occurs between watermark robustness, which aims to mitigate the potential for removal attacks discussed above, and spoofing—showing that robust watermarks are at an increased risk for spoofing attacks.
@@ 3 Preliminaries
Notations. Throughout the paper, we use x to denote a sequence of tokens, $\textbf{x}_{i}\in\mathcal{V}$ is the $i$-th token in the sequence, and $\mathcal{V}$ is the vocabulary. $M_{\text{orig}}$ denotes the original model without a watermark, $M_{\text{wm}}$ is the watermarked model, and $sk\in\mathcal{S}$ is the watermark secret key sampled from the key space $\mathcal{S}$.
Language Models. Current state-of-the-art LLMs are auto-regressive models, which predict the next token based on the prior tokens. Below, we more formally define language models (LM) considered in this work:
@@
We define a LM without a watermark as:
$M_{\text{orig}}:\mathcal{V}^{*}\rightarrow\mathcal{V},$ (1)
@@
Watermarks for LLMs. In this work, we focus on three SOTA decoding-based watermarking schemes: KGW [KGW+23a], Unigram [ZALW24] and Exp [KTHL23]. Informally, the decoding-based watermark is embedded by perturbing the output distribution of the original LLM. The perturbation is determined by secret watermark keys held by the LLM owner. More formally, we define the watermarking scheme as follows:
@@
$M_{\text{wm}}:\mathcal{V}^{*}\times\mathcal{S}\rightarrow\mathcal{V}$ (2)
The watermark detection outputs the statistical testing score for the null hypothesis that the input token sequence is independent of the watermark secret key:
$f_{\text{detection}}:\mathcal{V}^{*}\times\mathcal{S}\rightarrow\mathbb{R}$ (3)
The output score reflects the confidence of the watermark’s existence in the input. Please refer to Appendix A for the details of the three watermarks.
@@ 3.1 Threat Model
Attacker’s Objective. We study two types of attacks—watermark-removal attacks and spoofing attacks. In the watermark-removal attack, the attacker aims to generate a high-quality response from the LLM without an embedded watermark. For the spoofing attack, the goal is to generate a harmful or incorrect response that has the target victim organization’s watermark embedded.
Attacker’s Motivation. We present two practical scenarios to motivate watermark-removal attacks: 1) A student or a journalist uses high-quality watermarked LLMs to write articles, but they want to remove the watermark to claim originality. 2) A malicious company offering LLM services for clients, instead of developing their own LLMs, simply queries a watermarked LLM from a victim company and removes their watermarks, potentially infringing upon the intellectual property rights of the victim company.
In spoofing attacks, an attacker may wish to damage the reputation of a victim organization offering the LLM service. We again present two motivating examples: 1) The attacker can utilize the spoofing attack to generate fake news or incorrect facts and post them on social media. By claiming that fake or toxic material is generated by the benign company’s LLM, the attacker can add validity to their falsely generated content, and may damage the reputation of the company and their model. 2) The attacker could use the spoofing attack to inject malicious code into some public software. As the code has the benign company’s watermark embedded, and the benign company may be at fault and have to bear responsibility for the actions.
@@
Attacker’s Capabilities. In this work, we study attacks by exploiting three key properties of existing watermark schemes: 1) robustness, 2) quality preservation, and 3) publicly available detection APIs. The attackers have different capabilities and requirements for different attacks, but for each, we carefully limit our assumptions to those that are practical and easy to achieve in real-world deployment scenarios.
@@
1) For spoofing attacks exploiting robustness, we assume that the attacker can make $\mathsf{poly}(l)$ queries to the target watermarked LLM, where $l$ is the token length of the generated content. We also assume that the attacker can edit the generated sentence (e.g., insert or substitute tokens).
@@
2) For watermark-removal attacks exploiting quality preservation, we assume multiple watermark keys are utilized to embed the watermark, which is a common practice in designing robust cryptographic protocols and is suggested by SOTA watermarks [KGW+23a, KTHL23] to provide distortion-free guarantee in preserving the output quality or improve the security against brute-force attacks in distinguishing the watermarked tokens’ distribution. Similarly, for a sentence of length $l$, we assume the attacker can make $\mathsf{poly}(l)$ queries to the watermarked LLM under different watermark keys.
@@
3) For the scenarios of exploiting detection APIs, we assume that the detection API is available to normal users and the attacker can make $\mathsf{poly}(l)$ queries for a sentence of length $l$. The detection returns the watermark confidence score (p-value or z-score). For spoofing attacks exploiting the detection APIs, we assume that the attacker can auto-regressively synthesize (toxic) sentences. For example, they can run a local (small) model to synthesize such sentences. For watermark-removal attacks exploiting the detection APIs, we also assume that the attacker can make $\mathsf{poly}(l)$ queries to the watermarked LLM. As is common practice [NKIH23, OWJ+22] and also enabled by OpenAI’s API ($\mathrm{top\_logprobs=5}$), we assume that the top 5 tokens at each position and their probabilities are returned to the attackers.
@@ 4 Attacking Robust Watermarks
The goal of developing a watermark that is robust to output perturbations is to defend against watermark removal and has been the inspiration for recent works [ZALW24, KGW+23a, KTHL23, SKB+23, KGW+23b, PSF+23]. Formally, we define the watermark robustness:
@@
A watermark is $(\epsilon,\delta)$-robust, given a watermarked text x, all its neighboring texts within the $\epsilon$ editing distance, the probability that the detection fails to detect the edited text is bounded by $\delta$:
$\forall\textbf{x},\textbf{x}^{\prime}\in\mathcal{V}^{*},\,\Pr[f_{\text{detection}}(\textbf{x}^{\prime},sk)<T]<\delta,\quad s.t.\quad f_{\text{detection}}(\textbf{x},sk)\geq T,\,\text{d}(\textbf{x},\textbf{x}^{\prime})\leq\epsilon,$
where $T$ is the detection confidence threshold.
A more robust watermarking scheme can better defend against editing attacks aimed at removing the watermark. However, we note that this same property can be misused by malicious users to launch spoofing attacks—e.g., a small portion of toxic or incorrect content can be inserted into watermarked material, making it seem like it was generated by a specific watermarked LLM. With a robust watermark embedded, the entire toxic content will still seem watermarked, potentially damaging the reputation of the LLM service provider. We explore this more formally below.
Attack Procedure. 1) The attacker queries the target watermarked LLM to receive a high-entropy watermarked sentence $\textbf{x}_{\text{wm}}$, 2) The attacker edits $\textbf{x}_{\text{wm}}$ and forms a new piece of text $\textbf{x}^{\prime}$ and claims that $\textbf{x}^{\prime}$ is generated by the target LLM. The editing method can be defined by the attacker. Simple strategies could be inserting toxic tokens into the watermarked sentence $\textbf{x}_{\text{wm}}$ at random positions to make the whole content harmful or editing the tokens to make the sentence inaccurate (see the example in Table 1).
Attack Feasibility Analysis. We study the bound on the maximum number of tokens that are allowed to be inserted into a watermarked sentence, and we present the following theorem on Unigram due to its clean robustness guarantee:
Theorem 1 (Maximum insertion portion).
Consider a watermarked token sequence x of length $l$. The Unigram z-score threshold is $T$, the portion of the tokens in the green list is $\gamma$, the detection z-score of x is $z$, and the number of inserted tokens is $s$. Then to guarantee the expected z-score of the edited text is greater than $T$, it suffices to guarantee:
$\frac{s}{l}\leq\frac{z^{2}-T^{2}}{T^{2}}$ (4)
We present the proof of Theorem 1 in Appendix B. According to Eq. 4, as long as the number of inserted toxic tokens is bounded by $l\frac{z^{2}-T^{2}}{T^{2}}$, the attacker can execute a spoofing attack to generate toxic content with the target watermark embedded. For token insertion editing, the editing distance bound (Def. 3) for a sentence is $\epsilon=l\frac{z^{2}-T^{2}}{T^{2}}$. A stronger watermark increases the ease of launching spoofing attacks by allowing more toxic tokens to be inserted. This conclusion applies universally to all robust watermarking schemes. If a watermark is robust, such attacks are inevitable and may be extremely difficult to detect, as even one toxic token could potentially render the underlying content harmful or inaccurate (see, e.g., examples in Table 1).

### v3

# No Free Lunch in LLM Watermarking: Trade-offs in Watermarking Design Choices

@@ Abstract
Advances in generative models have made it possible for AI-generated text, code, and images to mirror human-generated content in many applications. Watermarking, a technique that aims to embed information in the output of a model to verify its source, is useful for mitigating the misuse of such AI-generated content. However, we show that common design choices in LLM watermarking schemes make the resulting systems surprisingly susceptible to attack—leading to fundamental trade-offs in robustness, utility, and usability. To navigate these trade-offs, we rigorously study a set of simple yet effective attacks on common watermarking systems, and propose guidelines and defenses for LLM watermarking in practice.
@@ 1 Introduction
Modern generative modeling systems have notably enhanced the quality of AI-produced content [4, 36, 29, 28]. For example, large language models (LLMs) like those powering ChatGPT [28] can generate text closely resembling human-crafted sentences. While this has led to exciting new applications of machine learning, there is also growing concern around the potential for misuse of these models, leading to a flurry of recent efforts on developing techniques to detect AI-generated content. A promising approach in this direction is to embed invisible watermarks into model-derived content, which can then be extracted and verified using a secret watermark key [17, 10, 6, 20, 46, 18, 13, 42, 39].
In this work, we identify that many of the key properties that make existing LLM watermarks successful can also render them susceptible to attack. In particular, we study a number of simple attacks that take advantage of common design choices of existing watermarking schemes, including:
@@
    Robustness of the watermarks to potential modifications in the output text, so that the watermarks cannot be easily removed [20, 17, 46, 20, 6, 13];
@@
    The use of multiple keys to prevent against watermark stealing attacks [17, 20, 15, 35, 11];
@@
    Public detection APIs, which allow the general public to easily verify whether or not candidate text is AI-generated [17, 37, 25].
While these common features and design choices of existing watermarking schemes have clear benefits, we show that they also make the resulting systems vulnerable to a number of simple but effective attacks. In particular, we study two types of attacks: 1) watermark-removal attacks, which remove the watermark from the watermarked content, and 2) spoofing attacks, which create (potentially toxic) content with a target watermark embedded, thus making the content appear to be generated by a specific LLM when it is in fact not (see motivating examples in Table 1).
@@
Z-Score: $5.98$uparrow
PPL: $4.46$
(a) Piggyback spoofing attack Alan Turing was born in 1950 and died in 1994, at the age of 43. He was the brilliant American scientist and mathematician who is largely credited with being the father of modern computer science. He is known for his contributions to mathematical biology and musicology. He was also one of the pioneers of computer science...
@@
Exploiting multiple keys
Z-Score: $2.40$downarrow
PPL: $4.05$
@@
Z-Score: $1.47$downarrow
PPL: $4.57$
Table 1: Examples generated using LLAMA-2-7B with/without the KGW watermark [17] under various attacks. We mark tokens in the green and red lists (see Appendix C). Z-score reflects the detection confidence of the watermark, and perplexity (PPL) measures text quality. (a) In the piggyback spoofing attack, we exploit watermark robustness by generating incorrect content that appears as watermarked (matching the z-score of the watermarked baseline), potentially damaging the reputation of the LLM. Incorrect tokens modified by the attacker are marked in orange and watermarked tokens in blue. (b-c) In watermark-removal attacks, attackers can effectively lower the z-score below the detection threshold while preserving a high sentence quality (low PPL) by exploiting either the (b) use of multiple keys or (c) publicly available watermark detection API.
Our work rigorously explores a number of simple removal and spoofing attacks for LLM watermarks. In doing so, we identify critical trade-offs that emerge between watermark robustness, utility, and usability as a result of watermarking design choices. To navigate these trade-offs, we propose potential defenses as well as a set of general guidelines to better enhance the security of next-generation LLM watermarking systems. Overall, we make the following contributions:
  • We study how watermark robustness, despite being a desirable property to mitigate removal attacks, can make the resulting systems highly susceptible to piggyback spoofing attacks, a simple type of attack that makes makes watermarked text toxic or inaccurate through small modifications, and show that challenges exist in detecting these attacks given that a single token can render an entire sentence inaccurate (Sec. 4).
  • We show that using multiple watermarking keys can make the system susceptible to watermark removal attacks (Sec. 5). Although a larger number of keys can help defend against watermark stealing attacks, which can be used to launch either spoofing or removal attacks, we show both theoretically and empirically that this in turn increases the potential for watermark removal attacks.
  • Finally, we identify that public watermark detection APIs can be exploited by attackers to launch both watermark-removal and spoofing attacks (Sec. 6). We propose a defense using techniques from differential privacy to effectively counteract spoofing attacks, showing that it is possible to avoid the possibilities of noise reduction by applying pseudorandom noise based on the input.
Throughout, we explore our attacks on three state-of-the-art watermarks [17, 46, 20] and two LLMs (LLAMA-2-7B [38] and OPT-1.3B [45])—demonstrating that these vulnerabilities are common to existing LLM watermarks, and providing caution for the field in deploying current solutions in practice without carefully considering the impact and trade-offs of watermarking design choices. Our code is available at https://github.com/Qi-Pang/LLM-Watermark-Attacks.
@@ 2 Related Work
Advances in large language models (LLMs) have given rise to increasing concerns that such models may be misused for purposes such as spreading misinformation, phishing, and academic cheating. In response, numerous recent works have proposed watermarking schemes as a tool for detecting LLM-generated text to mitigate potential misuse [17, 10, 6, 20, 46, 18, 13, 42, 39]. These approaches involve embedding invisible watermarks into the model-generated content, which can then be extracted and verified using a secret watermark key. Existing watermarking schemes share a few natural goals: (1) the watermark should be robust in that it cannot be easily removed; (2) the watermark should not be easily stolen, thus enabling spoofing or removal attacks; and (3) the presence of a watermark should be easy to detect when given new candidate text. Unfortunately, we show that existing methods that aim to achieve these goals can in turn enable simple but effective attacks.
Removal attacks. Several recent works have highlighted that paraphrasing methods may be used to evade the detection of AI-generated text [19, 14, 21, 22, 44], with [19, 44] demonstrating effective watermark removal using a local LLM. These methods usually require additional training for sentence paraphrasing which can impact sentence quality, or assume a high-quality oracle model to guarantee the output quality is preserved. In contrast, the simple and scalable removal attacks herein do not require additional training or a high-quality oracle. Additionally, our work differs in that we aim to directly connect and study how the inherent properties and design choices of watermarking schemes (such as the use of multiple keys and detection APIs) can inform such removal attacks.
Spoofing attacks. Prior works on spoofing use watermark stealing attacks to first estimate the watermark pattern and then embed it into an arbitrary content to launch spoofing attacks. These attacks usually require the attacker to pay a large startup cost by obtaining a significant number of watermarked tokens. For example, [35] requires 1 million queries to the watermarked LLM, and [15, 11] assume the attacker can obtain millions of watermarked tokens to estimate their distribution. Unlike these works, we explore spoofing attacks that are less flexible but can be launched with significantly less upfront cost. In Sec. 4, we explore a very simple and scalable form of spoofing exploiting the inherent robustness property of watermarks, which we refer to as a ‘piggyback spoofing attack’. In Sec. 6, we then explore more general spoofing attacks, which instead of querying the watermarked LLM numerous times, consider exploiting the public detection API. In both, our attacks do not require the attacker to estimate the watermark pattern, but share a similar ultimate goal with the prior spoofing attacks to create falsified inaccurate or toxic content that appears to be watermarked.
@@ 3 Preliminaries
Before exploring attacks and defenses on watermarking systems, we introduce relevant background on LLMs, notation we use throughout the work, and a set of concrete threat models.
Notation. We use x to denote a sequence of tokens, $\textbf{x}_{i}\in\mathcal{V}$ is the $i$-th token in the sequence, and $\mathcal{V}$ is the vocabulary. $M_{\text{orig}}$ denotes the original model without a watermark, $M_{\text{wm}}$ is the watermarked model, and $sk\in\mathcal{S}$ is the watermark secret key sampled from the key space $\mathcal{S}$.
Language Models. Current state-of-the-art (SOTA) LLMs are auto-regressive models, which predict the next token based on the prior tokens. We define language models more formally below:
@@
We define a language model (LM) without a watermark as:
$\footnotesize M_{\text{orig}}:\mathcal{V}^{*}\rightarrow\mathcal{V},$ (1)
@@
Watermarks for LLMs. In this work, we focus on three SOTA decoding-based watermarking schemes: KGW [17], Unigram [46] and Exp [20]. Informally, decoding-based watermarks are embedded by perturbing the output distribution of the original LLM. The perturbation is determined by secret watermark keys held by the LLM owner. Formally, we define the watermarking scheme:
@@
$\footnotesize M_{\text{wm}}:\mathcal{V}^{*}\times\mathcal{S}\rightarrow\mathcal{V}$ (2)
The watermark detection outputs the statistical testing score for the null hypothesis that the input token sequence is independent of the watermark secret key, which reflects the watermark confidence:
$\footnotesize f_{\text{detection}}:\mathcal{V}^{*}\times\mathcal{S}\rightarrow\mathbb{R}$ (3)
Please refer to Appendix C for additional details of the specific watermarks explored in this work.
@@ 3.1 Threat Model
Attacker’s Objective. We study two types of attacks—watermark-removal attacks and (piggyback or general) spoofing attacks. In the watermark-removal attack, the attacker aims to generate a high-quality response from the LLM without an embedded watermark. For the spoofing attacks, the goal is to generate a harmful or incorrect output that has the victim organization’s watermark embedded. We present concrete application scenarios for attacker’s motivations in Appendix B.
@@
Attacker’s Capabilities. We study attacks by exploiting three common design choices in watermarks: 1) robustness, 2) the use of multiple keys, and 3) public detection APIs. Each attack requires the adversary to have different capabilities, but we make assumptions that are practical and easy to achieve in real-world deployment scenarios.
@@
1) For piggyback spoofing attacks exploiting robustness (Sec. 4), we assume that the attacker can make $\mathcal{O}(1)$ queries to the target watermarked LLM. We also assume that the attacker can edit the generated sentence (e.g., insert or substitute tokens).
@@
2) For watermark-removal attacks exploiting the use of multiple keys (Sec. 5), we consider the scenario where multiple watermark keys are utilized to embed the watermark, which is a common practice in designing robust cryptographic protocols and is suggested by SOTA watermarks [20, 17] to improve resistance against watermark-stealing attacks [15, 11, 35]. For a sentence of length $l$, we assume that the attacker can make $\mathcal{O}(l)$ queries to the watermarked LLM.
@@
3) For the attacks on detection APIs (Sec. 6), we assume that the detection API is available to normal users and the attacker can make $\mathcal{O}(l)$ queries for a sentence of length $l$. The detection returns the watermark confidence score (p-value or z-score). For spoofing attacks exploiting the detection APIs, we assume that the attacker can auto-regressively synthesize (toxic) sentences. For example, they can run a local (small) model to synthesize such sentences. For watermark-removal attacks exploiting the detection APIs, we also assume that the attacker can make $\mathcal{O}(l)$ queries to the watermarked LLM. As is common practice [26, 32] and also enabled by OpenAI’s API [27], we assume that the top 5 tokens at each position and their probabilities are returned to the attackers.
@@ 4 Attacking Robust Watermarks
The goal of developing a watermark that is robust to output perturbations is to defend against watermark removal, which may be used to circumvent detection schemes for applications such as phishing or fake news generation. Robust watermark designs have been the topic of many recent works [46, 17, 20, 35, 18, 33]. We formally define watermark robustness in the following definition.
@@
A watermark is $(\epsilon,\delta)$-robust, given a watermarked text x, if for all its neighboring texts within the $\epsilon$ editing distance, the probability that the detection fails to detect the edited text is bounded by $\delta$, given the detection confidence threshold $T$:
$\footnotesize\forall\textbf{x},\textbf{x}^{\prime}\in\mathcal{V}^{*},\,\Pr[f_{\text{detection}}(\textbf{x}^{\prime},sk)<T]<\delta,\quad s.t.\,f_{\text{detection}}(\textbf{x},sk)\geq T,\,\text{d}(\textbf{x},\textbf{x}^{\prime})\leq\epsilon$
More robust watermarks can better defend against editing attacks, but this seemingly desirable property can also be easily misused by malicious users to launch simple piggyback spoofing attacks—e.g., a small portion of toxic or incorrect content can be inserted into the watermarked material, making it seem like it was generated by a specific watermarked LLM. The toxic content will still be detected as watermarked, potentially damaging the reputation of the LLM service provider. As discussed in Sec. 2, spoofing attacks explored in prior work usually require the attacker to obtain millions of watermarked tokens upfront to estimate the watermark pattern [15, 35, 11]. In contrast, our simple piggyback spoofing only requires a single query to the watermarked LLM with careful text modifications, and the effectiveness relates directly to the robustness of the LLM watermark.
Attack Procedure. (i) The attacker queries the target watermarked LLM to receive a high-entropy watermarked sentence $\textbf{x}_{\text{wm}}$, (ii) The attacker edits $\textbf{x}_{\text{wm}}$ and forms a new piece of text $\textbf{x}^{\prime}$ and claims that $\textbf{x}^{\prime}$ is generated by the target LLM. The editing method can be defined by the attacker. Simple strategies could include inserting toxic tokens into the watermarked sentence $\textbf{x}_{\text{wm}}$ at random positions, or editing specific tokens to make the output inaccurate (see example in Table 1). As we show, editing can also be done at scale by querying another LLM like GPT4 to generate fluent output.
We present the formal analysis on the attack feasibility in Appendix D and point out the takeaway that is universally applicable to all robust watermarks: A more robust watermark makes piggyback spoofing attack easier by allowing more toxic tokens to be inserted. This is a fundamental design trade-off: If a watermark is robust, such spoofing attacks are inevitable and may be extremely difficult to detect, as even one toxic token can render the entire content harmful or inaccurate.